# Data-driven Conditional Instrumental Variables for Debiasing Recommender Systems

### v1

@@ Introduction
where $z_{t_{u}}$ and $z_{c_{u}}$ indicate the representations of the CIV of the user $u$ and its conditional set, respectively, $z_{t_{p}}$ and $z_{c_{p}}$ denote the representations of CIV of the corresponding item $p$ and its conditional set, respectively.
Learning the Representations of CIV and Its Condition Set
In our CIV4Rec, we employ a VAE structure as our generative model for generating the representations of the CIV $Z_{t}$ and its conditional set $Z_{c}$ (Kingma and Welling 2013; Sohn, Lee, and Yan 2015) from the user interaction data $X$, referred to as the CIV and conditional Set Extraction Module (CSEM). We use the inference and generation networks of VAE to approximate the posterior distributions $p\left(Z_{t}|X\right)$ and $p\left(Z_{c}|X\right)$ for the two representations $Z_{t}$ and $Z_{c}$.
@@
The prior distribution $p\left(Z_{c}\right)$ is obtained based on the conditional VAE (CVAE) (Sohn, Lee, and Yan 2015) model. We use Monte Carlo (MC) sampling to condition $p(Z_{c})$ on $X$:
@@
$\begin{split}L_{civ}&=\mathbb{E}_{q}[\log p\left(X|Z_{t},Z_{c}\right)]\\ &-D_{KL}[q\left(Z_{t}|X\right)||p\left(Z_{t}\right)]\\ &-D_{KL}[q\left(Z_{c}|X\right)||p\left(Z_{c}|X\right)].\end{split}$ (9)
@@
Decomposition of Treatment Variable
After obtaining the representations of CIV ($Z_{t}$) and its conditional set ($Z_{c}$) from $X$, we use the CIV ($Z_{t}$) to reconstruct the treatment variable ($W$) and decompose $W$ to derive the causal relationship, specifically, user preference. We apply the least squares (LS) method based on IV4Rec (Si et al. 2022) to decompose $W$ and obtain the representation $\widehat{W}_{u,i}$ that is not affected by $U_{c}$:
@@
Click Prediction
Click prediction is the core task of a recommender system, with the goal of predicting user preferences. BPR loss is a commonly used optimization method for click prediction. We use BPR loss to optimize the reconstructed treatment variables to enhance click prediction:
$\begin{split}L_{click}=&-\sum_{\left(u,p,n\right)\in D}\ln sigma\big{(}\left<w_{u}^{re},w_{p}^{re}\right>\\ &-\left<w_{u}^{re},w_{n}^{re}\right>\big{)},\end{split}$ (14)
@@
Therefore, our CIV4Rec obtain $W_{u,i}^{re}$ for click prediction by learning the representations of CIV $Z_{t}$ and its condition set $Z_{c}$ from $X$ and by decomposing $W$.
Experiments
In this section, we conduct experiments on two real-world datasets to validate the recommendation performance and debiasing effectiveness of CIV4Rec.
Experimental Settings
We introduce the datasets, baselines, and experimental parameters used in the experiments. Detailed parameter settings are provided in the Appendix due to the page limit.
@@
We use two publicly available real-world datasets: the Movielens-10M dataset (Harper and Konstan 2015) and the Douban-Movie dataset [^1] https://moviedata.csuldw.com/. Both datasets include user IDs, movie IDs, and user ratings (1-5) for movies and are widely used in recommender system debiasing research. We binarized the datasets based on movie ratings, assigning a value of 1 to five-star ratings and 0 to the rest, and applied ten-core filtering to the data.
We follow the preprocessing steps used in DICE (Zheng et al. 2021) for the dataset. First, we randomly select 40% of the movie review data with the same probability (these data can be considered as unbiased and represent recommendation results under a completely random strategy). Then, we use 10% and 20% of the random data as the validation set and test set, respectively, to evaluate the model’s debiasing performance on unbiased data. Finally, the remaining 10% of the random data and the other 60% of unselected samples form the training set. In other words, the training, validation, and test data are split in a $70:10:20$ ratio. We include 10% unbiased data in the training set due to the requirements of CausE (Bonner and Vasile 2018). Table 1 shows the statistics of the processed datasets.
Dataset User # Item # Interaction #
@@
Causal debiasing methods are typically used as adjunct techniques alongside a backbone recommendation model. In our experiments, we use the mainstream recommendation models, MF and LightGCN, as the backbone. We compare our approach with the following five debiased recommendation methods based on causality:
  • IPS (Schnabel et al. 2016): This method assigns weights that are the inverse of an item’s popularity, thereby enhancing the impact of less popular items while reducing the influence of more popular ones.
  • IPS-C (Bottou et al. 2013): This approach caps the maximum value of IPS weights to reduce variance across the entire weight distribution.
  • CausE (Bonner and Vasile 2018): This method generates two sets of embeddings from the data, which are then aligned using regularization techniques to ensure their similarity.
  • DICE (Zheng et al. 2021): This method uses Structural Causal Modeling (SCM) (Pearl 2009) to define user-item interactions. This approach leverages the collision effect of causal reasoning to enhance training effectiveness.
  • DCCL (Zhao et al. 2023): This method uses contrastive learning to address data sparsity and the separation of these components.
@@
We evaluate Top-K recommendation performance on implicit feedback, a common setting in recommender systems. Top-K refers to the K items that the recommender system considers most relevant or attractive to the user, where K represents the number of items in the recommendation list. We use three frequently applied evaluation metrics: Recall, Hit Rate (HR), and NDCG, to evaluate all methods. The reported experimental results represent the optimal performance achieved by each method under its respective parameter settings.
It’s important to note that we did not compare our CIV4Rec method with IV-based debiasing methods like IV4Rec (Si et al. 2022) since these methods require specifying a specific in IV based on domain knowledge or expert nomination, such as self-collected user search data. In contrast, our CIV4Rec model is designed to learn the representations of CIV and its conditional set directly from user interaction data. The datasets used in our experiments are not suitable for IV-based debiasing methods.
Comparison of Experimental Results
@@
Tables 2 and 3 present the results of CIV4Rec and all baseline approaches on two real-world datasets. CIV4Rec-Causal is a variant of CIV4Rec that denotes click prediction using only the unbiased embeddings $\widehat{W}_{u,i}$, which capture the causal relationship. “Imp.” represents the percentage improvement in Recall for each method compared to the original model. The best results are highlighted in bold, and the second-best results are underlined.
@@
The analysis of Tables 2 and 3 shows that CIV4Rec and CIV4Rec-Causal significantly improve performance metrics compared to the original backbone, with the highest improvement reaching 39.71%, demonstrating statistical significance and the superiority of our approach. Notably, CIV4Rec consistently outperforms CIV4Rec-Causal, aligning with the understanding that incorporating appropriate confounders enhances recommendation performance. This confirms that $Z_{c}$ in CIV4Rec effectively captures relevant confounders in user interaction data.
@@
Tables 2 and 3 reveal several key insights: (1) IPS-based debiasing methods perform poorly because they rely heavily on the inverse propensity score, which is closely tied to the data distribution. In our experiments, training was conducted on a biased dataset, while testing was performed on an unbiased dataset, leading to inconsistent data distributions between the training and validation sets. (2) Similarly, CausE showed suboptimal performance because it requires an unbiased dataset during training to align the user-item embedding representations. (3) Although DICE and DCCL, two debiasing methods based on causal graph assumptions, improve recommendation performance, they do not achieve optimal results. This is because these methods target specific biases based on assumed causal graphs, while real-world datasets often contain diverse biases due to latent confounding factors, limiting their effectiveness.
Note that on the Movielens-10M dataset, while CIV4Rec’s improvement with LightGCN appears marginal compared to DICE, it is statistically significant, with a $p$-value of less than 0.05. This can be attributed to CIV4Rec’s ability to learn the CIV and its conditional set from interactive data, which reduces the typical constraints of such methods and effectively mitigates bias, leading to improved recommendation performance.
Evaluation on Debiasing Experiments Ability
@@
We use the Intersection Over Union (IOU(Zheng et al. 2021) metric to evaluate the debiasing ability of all methods. A higher IOU reflects more popular items in the recommendations, indicating weaker debiasing performance.
Figure 4 shows the IOU curves for all methods on the Douban-movie dataset. CIV4Rec and CIV4Rec-Causal exhibit the lowest IOU, indicating superior debiasing ability. Notably, the IOU for all baseline methods increases significantly as the number of recommended items grows, suggesting that their debiasing ability diminishes with more recommendations. In contrast, the debiasing ability of our CIV4Rec and CIV4Rec-Causal remains relatively stable, demonstrating greater robustness.
@@
Figure 4: The IOU of recommended items and popular items for all methods on the Douban-movie dataset. A larger IOU indicates a higher proportion of popular items in the recommendations, meaning the method has a weaker debiasing ability. (a) IOU of all methods on the MF; (b) IOU of all methods on the LightGCN.
@@
Figure 4 also illustrates that the IOU of CIV4Rec is higher than that of CIV4Rec-Causal, due to CIV4Rec incorporating confounding factor information. This suggests that $Z_{c}$ effectively captures confounding factors in user interaction data, validating our method.
Figure 5: Recall and IOU of MF-based CIV4Rec and its variants on the Douban-movie dataset.
Ablation Studies
@@
We perform ablation studies to evaluate the effectiveness of each component in CIV4Rec. To verify the effectiveness of CIV and its condition set, we propose CIV4Rec-Con, which uses only $Z_{c}$ for click prediction. Figure 5 presents the IOU and Recall of CIV4Rec and its variants. The results show that CIV4Rec-Con has the highest IOU and lowest Recall, highlighting the effectiveness of $Z_{c}$ in capturing confounding factor information. In contrast, CIV4Rec-Causal exhibits higher Recall but lower IOU than CIV4Rec-Con, indicating its effectiveness in capturing user preference information and mitigating confounding factors through $Z_{t}$. CIV4Rec, by integrating both user preference and confounding factor information, achieves higher Recall and IOU than CIV4Rec-Causal, demonstrating the combined effectiveness of $Z_{t}$ and $Z_{c}$.
Conclusion
@@
In this paper, we propose a data-driven CIV debiasing method called CIV4Rec. We learn the representations of CIV and its conditional set from user interaction data. The CIV is used to decompose the treatment variable and uncover the causal relationships between variables, while the conditional set captures confounding factors in the user interaction data. Unlike existing IV-based debiasing methods, CIV4Rec imposes fewer constraints and does not require the selection of specific IVs based on domain knowledge. By integrating confounding factors and the causal relationships of the treatment variable, CIV4Rec achieves high-quality recommendations and effective debiasing. We conducted extensive experiments on two real-world datasets to validate the effectiveness and superiority of CIV4Rec in both recommendation and debiasing performance.
Appendix A Preliminaries
In this section, we will introduce the fundamental concepts of causal inference related to our main manuscript.
Directed Acyclic Graph
Causal graphs use Directed Acyclic Graphs (DAGs) to represent causal relationships between variables. In these graphs, nodes represent variables, and edges represent their relationships. Specifically, there are three classical DAGs used to describe causal relationships: the chain $A\rightarrow B\rightarrow C$, the fork $A\leftarrow B\rightarrow C$, and the collider $A\rightarrow B\leftarrow C$.
In the chain DAG $A\rightarrow B\rightarrow C$, $A$ influences $C$ through the intermediary $B$. In the fork DAG $A\leftarrow B\rightarrow C$, $B$ is the confounder or common cause of $A$ and $C$, meaning $B$ influences both $A$ and $C$, resulting in a correlation between $A$ and $C$. However, this correlation does not imply a direct causal relationship between $A$ and $C$. The collider structure $A\rightarrow B\leftarrow C$ describes a situation where $A$ and $C$ are independent of each other but jointly influence the collision node $B$. In this case, $A$ and $C$ correlate when conditioned on $B$.
D-Separation
$D$-separation is a key concept in graphical models (Pearl 2009). It helps determine if a set of variables is conditionally independent given another set of variables within a DAG. It’s an essential graph criterion for understanding conditional dependence or independence in a DAG.
Definition 2 ($d$-Separated (Pearl 2009)).
A path is said to be d-separated (or blocked) by conditioning on a set of nodes $Z$ if and only if one of the two conditions is satisfied:
  1. 1.
    The $path$ contains a chain structure or a fork structure, and the middle node $B$ belongs to $Z$.
  2. 2.
    The $path$ contains a collision structures and the collision node $B$ and its descendant nodes are not in $Z$.
where a $path$ represents a sequence of consecutive edges (directionless) in a DAG. For example, in causal DAGs $A\rightarrow B\rightarrow C$ and $A\leftarrow B\rightarrow C$, $A$ and $C$ are d-separated by conditioning on $B$, whereas in $A\rightarrow B\leftarrow C$, $A$ and $C$ are d-separated when $B$ is not conditioned upon.
Conditional Instrumental Variable
The Instrumental Variable (IV) approach is powerful for addressing confounding factors in causal inference. As the Introduction mentions, the last two conditions of a standard IV (i.e., IV must be independent of confounders and outcome variables) are often too strict for real-world applications. In contrast, a conditional IV (CIV) has more relaxed conditions than a standard IV (Brito and Pearl 2012; Pearl 2009).
Definition 3 (Conditional IV (Brito and Pearl 2012; Pearl 2009)).
Given a DAG $\mathcal{G}=\left(V,E\right)$, where nodes $V=X\cup U_{C}\cup\{W,Y\}$ and $E$ are directed edges. $W$ and $Y$ are the treatment and outcome variables, respectively. $X$ is the set of measured pretreatment variables, and $U_{C}$ is the unmeasured confounders. $Z_{c}\subseteq X\backslash\{Z_{t}\}$ and $Z_{c}\neq\emptyset$. A variable $Z_{t}$ is a CIV if it satisfies the following three conditions:
  1. 1.
    $Z_{c}$ contains only non-descendants of $Y$;
  2. 2.
    When conditioning on $Z_{c}$, $Z_{t}$ and $W$ are d-connection;
  3. 3.
    When conditioning on $Z_{c}$, $Z_{t}$ and $Y$ are d-separation in the manipulated graph $\mathcal{G}_{\underline{W}}$, which is obtained by deleting all arrows emerging from nodes in $W$ from $\mathcal{G}$.
Figure 6: A causal DAG $\mathcal{G}$ shows the CIV exemplar.
As an example of the application of this definition, Figure 6 shows the causal DAG of CIV. In the manipulated graph $\mathcal{G}_{\underline{W}}$, after conditioning on variable $Z_{c}$, $Z_{t}$ and $Y$ are d-separation but d-connection with $W$, and $Z_{c}$ contains only non-descendants of $Y$.
Appendix B The CIV and Its Condition Set Extraction Module
Figure 7 shows the architecture of our latent representation of CIV $Z_{t}$ and its conditional set $Z_{c}$ from user interaction data $X$. We use the inference and generation networks of VAE (Kingma and Welling 2013) to approximate the posterior distributions $p\left(Z_{t}|X\right)$ and $p\left(Z_{c}|X\right)$ of $Z_{t}$ and $Z_{c}$.
Appendix C Parameter Settings
To ensure fair comparisons, we standardised the parameter counts across all methods. We set the embedding size for models utilising DICE (Zheng et al. 2021) and DCCL (Zhao et al. 2023) to 64 since they comprise two concatenated sets of embeddings. For the other models, we maintained a consistent embedding size of 128. In the CIV4Rec method, the hyperparameter $\alpha$ was set to 0.85. We employed the Adam Optimiser to update model weights, with an initial learning rate of 0.001 and a batch size 128. Except for DICE and DCCL, the loss functions for the other baseline models were based on the BPR (Rendle et al. 2012) function. All models were executed on an NVIDIA A100 (40GB RAM) GPU. To assess the model performance of our approach, we utilised three widely recognised metrics in the recommender systems field: Recall, Hit Ratio (HR), and NDCG.

### v2

@@ 3.3 The Concepts of Treatment Variable, CIV and Its Conditional Set
where $z_{t_{u}}$ and $z_{c_{u}}$ indicate the representations of the CIV of the user $u$ and its conditional set, respectively, $z_{t_{p}}$ and $z_{c_{p}}$ denote the representations of CIV of the corresponding item $p$ and its conditional set, respectively. To estimate the causal effects of users and items on $Y$ (user feedback), it is necessary to construct the CIV and its conditional set for the user $u$ and item $p$, respectively.
3.4 Learning the Representations of CIV and Its Condition Set
In our IDCIV-RS framework, we employ a VAE structure as the generative model to generate the representations of the CIV $Z_{t}$ and its conditional set $Z_{c}$ Kingma and Welling (2013); Sohn et al. (2015) from the user interaction data $X$, referred to as the CIV and conditional Set Extraction Module (CSEM), as shown in Figure 4. We use the inference and generation networks of VAE to approximate the posterior distributions $p\left(Z_{t}|X\right)$ and $p\left(Z_{c}|X\right)$ for the two representations $Z_{t}$ and $Z_{c}$.
@@ 3.4 Learning the Representations of CIV and Its Condition Set
The prior distribution $p\left(Z_{c}\right)$ is obtained based on the Conditional VAE (CVAE) Sohn et al. (2015) model. We use Monte Carlo (MC) sampling to condition $p(Z_{c})$ on $X$:
@@
$\begin{split}L_{civ}=&\mathbb{E}_{q}[\log p\left(X|Z_{t},Z_{c}\right)]-D_{KL}[q\left(Z_{t}|X\right)||p\left(Z_{t}\right)]\\ &-D_{KL}[q\left(Z_{c}|X\right)||p\left(Z_{c}|X\right)].\end{split}$ (9)
@@
3.5 Decomposition of Treatment Variable
After obtaining the representations of CIV ($Z_{t}$) and its conditional set ($Z_{c}$) from $X$, we use the CIV ($Z_{t}$) to reconstruct the treatment variable ($W$) and decompose $W$ to derive the causal relationship, specifically, user preference. We apply the least squares (LS) method based on IV4Rec Si et al. (2022) to decompose $W$ and obtain the representation $\widehat{W}_{u,i}$ that is not affected by $U_{c}$:
@@ 3.5 Decomposition of Treatment Variable
3.6 Click Prediction
To improve click prediction, we optimize the reconstructed treatment variables with BPR loss.
@@ 3.6 Click Prediction
where $\left$$\cdot\right$> denotes the inner product, and $w_{n}$ is the negative sample item that user $u$ has not interacted with, selected by the Popularity-based Negative Sampling with Margin (PNSM) strategy Zheng et al. (2021). By combining the ELBO and BPR, the final loss function of our IDCIV-RS is:
@@
Therefore, our IDCIV-RS obtain $W_{u,i}^{re}$ for click prediction by learning the representations of CIV $Z_{t}$ and its condition set $Z_{c}$ from $X$ and by decomposing $W$.
4 Experiments
In this section, we conduct experiments on two real-world datasets to validate the recommendation performance and debiasing effectiveness of IDCIV-RS.
4.1 Experimental Settings
@@
We use two publicly real-world datasets: the Movielens-10M and the Douban-Movie datasets. Both datasets include user IDs, movie IDs, and ratings (1-5) for movies and are widely used in recommender system debiasing research. We use a dataset setup commonly used in the field of debiased recommender systems; specifically, the training set is biased and the test set is unbiased. We adopt the data pre-processing approach used in previous studies Zheng et al. (2021).
Dataset # User # Item # Interaction
@@
Causal debiasing methods are typically applied as enhancements to backbone recommendation models. In our experiments, we use MF and LightGCN as the backbone models. We compare our approach against five causality-based debiasing methods:
  • IPS Schnabel et al. (2016): This method assigns weights that are the inverse of an item’s popularity, thereby enhancing the impact of less popular items while reducing the influence of more popular ones.
  • IPS-C Bottou et al. (2013): This approach caps the maximum value of IPS weights to reduce variance across the entire weight distribution.
  • CausE Bonner and Vasile (2018): This method generates two sets of embeddings from the data, which are then aligned using regularization techniques to ensure their similarity.
  • DICE Zheng et al. (2021): This method uses Structural Causal Modeling (SCM) Pearl (2009) to define user-item interactions. This approach leverages the collision effect of causal reasoning to enhance training effectiveness.
  • DCCL Zhao et al. (2023): This method uses contrastive learning to address data sparsity and the separation of these components.
We did not compare IDCIV-RS with IV-based methods like IV4Rec Si et al. (2022), as these require explicit IVs derived from domain knowledge or user search data, which our datasets lack. In contrast, IDCIV-RS learns CIV representations directly from user interactions, avoiding dependence on unavailable or domain-specific data. This confers greater flexibility and applicability in real-world scenarios.
@@
We evaluate Top-K recommendation under implicit feedback using Recall, Hit Rate (HR), and NDCG. Results reflect each method’s best performance under optimal settings. “Imp.” denotes the percentage improvement in Recall over the base model.
4.2 Comparison of Experimental Results
@@
Tables 2 and 3 present the results of IDCIV-RS and all baseline approaches on two real-world datasets. IDCIV-RS-Causal is a variant of IDCIV-RS that denotes click prediction using only the unbiased embeddings $\widehat{W}_{u,i}$, which capture the causal relationship.
@@
The analysis of Tables 2 and 3 shows that IDCIV-RS and IDCIV-RS-Causal significantly improve performance metrics compared to the original backbone, with the highest improvement reaching 39.71%, demonstrating statistical significance and the superiority of our approach. Notably, IDCIV-RS consistently outperforms IDCIV-RS-Causal, aligning with the understanding that incorporating appropriate confounders enhances recommendation performance. This confirms that $Z_{c}$ in IDCIV-RS effectively captures relevant confounders in user interaction data.
@@
Tables 2 and 3 provide several key insights: (1) IPS-based debiasing methods perform poorly due to their reliance on the inverse propensity score, which is sensitive to data distribution. In our experiments, training on a biased dataset and testing on an unbiased one led to distribution mismatches. (2) CausE also underperforms, as it requires an unbiased training dataset to align user-item embeddings. (3) Although DICE and DCCL, which are based on causal graph assumptions, improve performance, they still fall short of optimal results. This is because they target specific biases based on predefined causal graphs, while real-world datasets often contain diverse biases from latent confounders, limiting their effectiveness.
4.3 Evaluation on Debiasing Experiments Ability
@@
We use the Intersection Over Union (IOUZheng et al. (2021) metric to evaluate the debiasing ability of all methods. A higher IOU reflects more popular items in the recommendations, indicating weaker debiasing performance.
@@
Figure 5: The IOU of recommended items and popular items for all methods on the Douban-movie dataset. (a) IOU of all methods on the MF; (b) IOU of all methods on the LightGCN.
Figure 5 shows the IOU for all methods on the Douban-movie dataset. IDCIV-RS and IDCIV-RS-Causal exhibit the lowest IOU, indicating superior debiasing ability. Notably, the IOU for all baseline methods increases significantly as the number of recommended items grows, suggesting that their debiasing ability diminishes with more recommendations. In contrast, the debiasing ability of our IDCIV-RS and IDCIV-RS-Causal remains relatively stable, demonstrating greater robustness.
@@
Figure 5 also illustrates that the IOU of IDCIV-RS is higher than that of IDCIV-RS-Causal, due to IDCIV-RS incorporating confounding factor information. This suggests that $Z_{c}$ effectively captures confounding factors in user interaction data, validating our method.
4.4 Ablation Studies
@@
We perform ablation studies to evaluate the effectiveness of each component in IDCIV-RS. To verify the effectiveness of CIV and its condition set, we propose IDCIV-RS-Con, which uses only $Z_{c}$ for click prediction. Figure 6 presents the IOU and Recall of IDCIV-RS and its variants. The results show that IDCIV-RS-Con has the highest IOU and lowest Recall, highlighting the effectiveness of $Z_{c}$ in capturing confounding factor information. In contrast, IDCIV-RS-Causal exhibits higher Recall but lower IOU than IDCIV-RS-Con, indicating its effectiveness in capturing user preference information and mitigating confounding factors through $Z_{t}$. IDCIV-RS, by integrating both user preference and confounding factor information, achieves higher Recall and IOU than IDCIV-RS-Causal, demonstrating the combined effectiveness of $Z_{t}$ and $Z_{c}$.
Figure 6: Recall and IOU of MF-based IDCIV-RS and its variants on the Douban-movie dataset. Where the bar represents Recall and the curve represents IOU.
5 Conclusion
@@
In this paper, we propose a data-driven CIV debiasing method called IDCIV-RS. We learn the representations of CIV and its conditional set from user interaction data. The CIV is used to decompose the treatment variable and uncover the causal relationships between variables, while the conditional set captures confounding factors in the user interaction data. Unlike existing IV-based debiasing methods, IDCIV-RS imposes fewer constraints and does not require the selection of specific IVs based on domain knowledge. By integrating confounding factors and the causal relationships of the treatment variable, IDCIV-RS achieves high-quality recommendations and effective debiasing. We conducted extensive experiments on two real-world datasets to validate the effectiveness and superiority of IDCIV-RS in both recommendation and debiasing performance.
Acknowledgments
This work is supported partly by the Project of Guangxi Science and Technology (2025GXNSFFA069015), the National Natural Science Foundation of China (No. 62372119, 62166003), the Key Lab of Education Blockchain and Intelligent Technology, Ministry of Education (No. EBME24-03), the Research Fund of the Guangxi Key Lab of Multi-source Information Mining & Security (MIMS24-M-01), and the Australian Research Council (under grant DP230101122).